Test Beam Shift Bulletin Board¶
Welcome to the NOvA Test Beam shift bulletin board! Look here for relevant information about the current status of operations.
Run Coordinating & Experts¶
- Run Coordinator: Mike Wallbank (331-814-6766), Yagmur Torun (312-420-5519). (Run Coordinator Schedule).
- Deputy Run Coordinator: Andrew Sutton (919-619-8866)
- See here for current NOvA DAQ on-call experts: http://dbweb6.fnal.gov:8080/ECL/nova/C/show_month
- See Troubleshooting on Shift for solutions to common problems when on shift.
- Problems with the beamline DAQ or DQM? Trying restarting both before contacting an expert.
- The Beamline VNC (VNC4) is now hosted on the novabeamlinedaq00 machines. You no longer need to ssh to that machine to get onto it (to, e.g. run the DAQ, launch a new msgviewer etc).
- The 'Start Beamline Run' and 'End Beamline Run' forms are now automated by the DAQ, so the shifter no longer needs to fill these out.
- Use the 'Beamline info' icon on VNC 4 to get all relevant information about the beamline configuration.
- We have switched to using the non-expert version of Run Control for the detector DAQ. What does this mean for you? Fewer button clicks to start a run! You should just need to click on
Configure Resources, and
Start Run. It will automatically pop-up the usual list of resources for you to okay, and the window for you to select the configuration.
Current Running Conditions¶
- Run both detector and beamline DAQs to collect data.
Known and Upcoming Downtimes¶
- Friday, March 20, time TBD -- Tuesday, March 24, 4pm
- NOTE: No access to WH at all March 21-March 22, 6:30am-4:30pm
Current DAQ Configurations¶
on: PADDLE1 PADDLE2 PADDLE3 PADDLE4 TRIGCLEAR SWYDGATE
- Configuration: TestBeam_FebV4_11_33_test
- We still have an issue with triggers dropping off after we've been running for awhile (this occurs ~1-2 times a day). If you see the "DDT: Activity1" rate drop below 10 Hz, please restart the detector run. Other indications of the issue include the EVD updating less frequency than normal, DCM activity dropping off in Ganglia, and no beamline triggers during a spill (this last one could just be a sign that beam is down, so don't take it as evidence by itself).
- Recently (since January 16th), we have seen instances of the TDU disconnecting in TDUControl. When the TDU disconnects the indicator next to
GPS Lockwill turn red. If this happens, please post a screenshot of TDUControl in the ECL, and then hit the
Connectbutton to reconnect. Keep a close eye on FEB shutoffs after this to make sure the re-enable DAQ is working properly.
- The detector is now interlocked with environmental monitoring, so if there is an environment alarm, the HV and LV power supplies will shutdown, preventing damage to APDs and FEBs. However, you should keep an eye on the "PLC Dry Gas & Water" synoptic display, in case the HVAC freezes or the water chiller trips. If the HVAC stops cooling MC7 (most likely triggering an environment dew point alarm), the water chiller should be stopped, to prevent condensation on the water lines flowing through the detector, until the HVAC is recovered. If the water chiller stops working (most likely triggering an Inlet APD Water Temp alarm), a picture of the front panel in the water chiller should be taken, and the water chiller should be stopped until it can be recovered. More information is found under: Environment Emergency Procedures
- If the VNC crashes, it can be reopened from the icon on the desktop. This will pop up a terminal and open up the display. If the terminal doesn't close on its own, please close it yourself otherwise the VNC may not start properly. If the icon does not work, Get a Kerberos Ticket and try again.